Versions Compared


  • This line was added.
  • This line was removed.
  • Formatting was changed.

11:00 - 11:30   Replaced the tsoil probe and PIC at tnw05. Data looks good now.

UPDATE:  Tsoil09 removed from this site had its eeprom space cleared. The probe is now working once the eeprom space was restored.